# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ Appendix B Photon counting and calculation of photon generation probabilities
Photon detection and counting lies at the heart of Quantum Optics and thus plenty of methods have been developed over time Gardiner and Zoller (2004); Plenio and Knight (1998). In this Appendix we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs. In Sec. B.1 we introduce an original photon counting method that we developed based on extending the master equation formalism. Then, in Sec. B.2, we discuss a more standard photon counting method using the quantum trajectory (QT) approach. This is useful when the Hilbert space of the system becomes very large, but at the expense of losing precision in the computation of the averages.
@@ B.1 Photon counting in master equation formalism
In the master equation dynamics the information about the emission of photons into the bath is omitted, and therefore one typically resorts to the input-output formalism (cf. Sec. C and Ref. Gardiner and Zoller (2004)) in order to relate measurable photonic quantities to multi-time system correlations. Nevertheless, obtaining the photon statistics from these system correlations involves the computation of multi-dimensional integrals over time, which can be very computationally costly and inefficient when scaling up the number of emitters or photons to probe.
@@
but as long as the total dimension fits in ${\rm dim}\{{\cal H}_{\rm ext}\}\lesssim 2^{10}$ our method provides a fast and efficient way to numerically calculate the few-photon statistics of multiple emitters within a purely master equation approach. For any Markovian master equation for a system state $\tilde{\rho}(t)$, one can find an extended master equation that incorporates the few-photon counting statistics in the dynamics. The general recipe is actually very simple and thus we explain it directly on the synchronized multiphoton device described by Eq. (12). In this case, the extended master equation reads
@@
In practical calculations we therefore solve for the extended state $\rho_{\rm ext}(t)$ in Eq. (54), and we then obtain the few-photon statistics of the emitted photons by taking simple expectation values on $\rho_{\rm ext}(t)$. In particular, the probability ${\cal P}_{n}^{j}$ to count $n$ photons in channel $j$ is calculated as
@@
In the general case, however, it is important to ensure that the occupation of the last state of the counters is negligible, so that the photon statistics is not affected by the finite size of the counters.
@@ B.2 Photon counting in quantum trajectories formalism
The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al. (1992); Mølmer et al. (1993); Dum et al. (1992). Here, the physics of quantum jumps is explicitly simulated during the open system evolution and therefore it is very natural to count them and thereby infer the photon statistics.
@@
When calculating many trajectories $M\gg 1$ of the system dynamics, we can gather enough quantum jump data and determine the photon emission probabilities (71) and (72) with low statistical error. Since this error decreases as $\sim M^{-1/2}$, we typically require on the order of $M\sim 10^{3}-10^{4}$ trajectories to obtain meaningful results with an error on order $\sim 10^{-2}-10^{-3}$. In practice, this makes the calculation of ${\cal P}^{j}_{n}$ and $P_{N}$ less precise than the extended master equation method in Sec. B.1. Nevertheless, the advantage of QT formalism is that we evolve pure states instead of density matrices and that it avoids extending the Hilbert space dimension to include counters as in the extended master equation method. This two key aspects allows us to dramatically reduce the Hilbert space for the simulations and thus to treat much larger systems composed of many more SPSs [cf. the scalability calculations in Fig. 6].
@@ Appendix C Photon correlations and input-output formalism
In Secs. III.4-III.5, we discuss how to quantify multi-photon contamination and photon indistinguishability in the emission of SPSs via second-order photon correlation functions. These photon correlations are measured via coincidence counts either in the Hanbury Brown and Twiss (HBT) or the Hong-Ou-Mandel (HOM) configurations, and can be expressed as

### v3

@@ Appendix B Photon counting and calculation of photon generation probabilities
Photon detection and counting lies at the heart of Quantum Optics and thus plenty of methods have been developed over time Gardiner and Zoller (2004); Plenio and Knight (1998). In this Appendix, we describe two methods for quantifying the efficiencies and photon statistics from the output of many SPSs. In Sec. B.1 we introduce an original photon counting method that we developed based on extending the master equation formalism. Then, in Sec. B.2, we discuss a more standard photon counting method using the quantum trajectory (QT) approach. This is useful when the Hilbert space of the system becomes very large but at the expense of losing precision in the computation of the averages.
@@ B.1 Photon counting in master equation formalism
In the master equation dynamics, the information about the emission of photons into the bath is omitted, and therefore one typically resorts to the input-output formalism (cf. Sec. C and Ref. Gardiner and Zoller (2004)) to relate measurable photonic quantities to multi-time system correlations. Nevertheless, obtaining the photon statistics from these system correlations involves the computation of multi-dimensional integrals over time, which can be very computationally costly and inefficient when scaling up the number of emitters or photons to probe.
@@
but as long as the total dimension fits in ${\rm dim}\{{\cal H}_{\rm ext}\}\lesssim 2^{10}$ our method provides a fast and efficient way to numerically calculate the few-photon statistics of multiple emitters within a purely master equation approach. For any Markovian master equation for a system state $\tilde{\rho}(t)$, one can find an extended master equation that incorporates the few-photon counting statistics in the dynamics. The general recipe is very simple and thus we explain it directly on the synchronized multiphoton device described by Eq. (13). In this case, the extended master equation reads
@@
In practical calculations, we therefore solve for the extended state $\rho_{\rm ext}(t)$ in Eq. (63), and we then obtain the few-photon statistics of the emitted photons by taking simple expectation values on $\rho_{\rm ext}(t)$. In particular, the probability ${\cal P}_{n}^{j}$ to count $n$ photons in channel $j$ is calculated as
@@
In the general case, however, it is important to ensure that the occupation of the last state of the counters is negligible so that the photon statistics are not affected by the finite size of the counters.
@@ B.2 Photon counting in quantum trajectories formalism
The most natural way to implement an ideal photon counting is within the formalism of quantum trajectories (QT) and continuous measurements Plenio and Knight (1998); Daley (2014); Dalibard et al. (1992); Molmer et al. (1993); Dum et al. (1992). Here, the physics of quantum jumps is explicitly simulated during the open system evolution and therefore it is very natural to count them and thereby infer the photon statistics.
@@
When calculating many trajectories $M\gg 1$ of the system dynamics, we can gather enough quantum jump data and determine the photon emission probabilities (80) and (81) with low statistical error. Since this error decreases as $\sim M^{-1/2}$, we typically require on the order of $M\sim 10^{3}-10^{4}$ trajectories to obtain meaningful results with an error on order $\sim 10^{-2}-10^{-3}$. In practice, this makes the calculation of ${\cal P}^{j}_{n}$ and $P_{N}$ less precise than the extended master equation method in Sec. B.1. Nevertheless, the advantage of QT formalism is that we evolve pure states instead of density matrices and that it avoids extending the Hilbert space dimension to include counters as in the extended master equation method. These two key aspects allow us to dramatically reduce the Hilbert space for the simulations and thus to treat much larger systems composed of many more SPSs [cf. the scalability calculations in Fig. 6].
@@ Appendix C Photon correlations and input-output formalism
In Secs. III.4-III.5, we discuss how to quantify multi-photon contamination and photon indistinguishability in the emission of SPSs via second-order photon correlation functions. These photon correlations are measured via coincidence counts either in the Hanbury Brown and Twiss (HBT) or the Hong-Ou-Mandel (HOM) configurations and can be expressed as